OpenStudy (starwars18):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation.

OpenStudy (starwars18):

\[\frac{ \sin x }{ 1 - \cos x } + \frac{ \sin x }{ 1 + \cos x } = 2 \csc x\]

OpenStudy (dan815):

okay first thing to know is csc(x) = 1/sin(x)

OpenStudy (starwars18):

yes

OpenStudy (dan815):

you should simplify the left side, by multiplying both fracctions by the conjugate

OpenStudy (dan815):

1-cos(x) and 1+cos(x) are factors of 1-(cos(x))^2

OpenStudy (dan815):

and 1-cos^2(x) = sin^2(x) as sin^2(x) + cos^2(x) = 1 this identity comes from the unit circle

OpenStudy (dan815):

okay yep it is right 2/sin(x) = 2csc(x) should be the right answer

OpenStudy (starwars18):

so sin x * (1 + cos x) + sin^3 x

OpenStudy (starwars18):

so they all cancel out leaving 2 over sin x... i get it now

OpenStudy (anonymous):

2 / sin(x) is like saying 2 times 1 /sin(x) which mean 2 times csc(x) or 2 csc(x) since csc = 1/sin(x)

OpenStudy (starwars18):

yes, ty i got it. sorry i forgot to close the question!!!